OpenStudy (anonymous):

Find an equation in standard form of the line having the given slope and and containing the given point. m=-6, (5,6) a.) 6x+y=36 b.) -6x-y=41 c.)-6x+y=41 d.) 6x-y=36

OpenStudy (bahrom7893):

y-6=-6(x-5) y-6=-6x+30 y+6x=36

OpenStudy (bahrom7893):

or 6x+y=36